OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions.Check all that apply. 8y - 24x = 64 9y + 45x = 72 inconsistent infinitely many solutions consistent dependent one solution independent

OpenStudy (anonymous):

Is this a linear algebra question by chance?

OpenStudy (darkbluechocobo):

Yes this is all about graphing ._.

OpenStudy (darkbluechocobo):

I realize with this one the lines intersect at (0,8) both of them

OpenStudy (darkbluechocobo):

And actually i take that back this is Algebra 2

OpenStudy (anonymous):

Ok cuz i was going to use some linear algebra techniques but nevermind lol. But from your graph you can see that the system is consistent, has one solution, and independent from each other. You can tell they're independent because the lines are not scalar multiples of each other, in other words you cannot multiply one equation by a number and get the second equation.

OpenStudy (darkbluechocobo):

O: Thank you good sir Thank you

OpenStudy (darkbluechocobo):

Can you actually help me out with one more Then I am done

OpenStudy (anonymous):

Sure

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions. Check all that apply. $2x + 2y = -10$ $4x + 4y = -16$ consistent infinitely many solutions inconsistent no solution

OpenStudy (darkbluechocobo):

stupid dollar signs ._.

OpenStudy (darkbluechocobo):

Imagine the dollar signs aren't there

OpenStudy (darkbluechocobo):

These are parallel aren't they o.o

OpenStudy (anonymous):

alright, if you divide the second equation by 2 you can see that the lines have the same slope, meaning they are parallel. Parallel lines never cross eachother, therefore inconsistent and no solution

OpenStudy (darkbluechocobo):

and no solution o.o ??

OpenStudy (anonymous):

yes, inconsistent is a fancy way of saying no solution so they both are true.

OpenStudy (darkbluechocobo):

*facepalm* Thank you again sir. You have a new fan :D

OpenStudy (anonymous):

Good deal! Glad i could help

OpenStudy (darkbluechocobo):

:D Now if you ever need help I shall try if possible

OpenStudy (anonymous):

Sounds like a plan!

OpenStudy (darkbluechocobo):

And actually could you check me on these 2 problems I am unsure about

OpenStudy (darkbluechocobo):

Sorry but its always safe to know whether right or wrong v.v

OpenStudy (anonymous):

sure lets see them.

OpenStudy (darkbluechocobo):

Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -3). $3y + 6x = 3$ $x - y = -7$ Well I said No to this one

OpenStudy (anonymous):

is the question the same as the two ive helped with before?

OpenStudy (darkbluechocobo):

No, Basically you fill in the ordered pair with x and y

OpenStudy (anonymous):

or it wants a solution is what it sounds like

OpenStudy (darkbluechocobo):

so like 3(-3) +6(5)

OpenStudy (anonymous):

alrighty it wants us to use a graph, so we gotta solve these for y so we can graph them

OpenStudy (darkbluechocobo):

Oks

OpenStudy (anonymous):

1. y = -2x + 1 2. y = x + 7

OpenStudy (darkbluechocobo):

it intersects at like -3.5

OpenStudy (anonymous):

looks like (-2,5)

OpenStudy (darkbluechocobo):

what does that mean o.o

OpenStudy (darkbluechocobo):

Wait so that means yes

OpenStudy (anonymous):

thats the point that they intersect at. But to be sure plug in the point (-2,5) to see if it satisfies both equations so for the 1st one \[5 = -2(-2) + 1\]\[5 = 4 + 1 \] that one works now for the second one \[5 = -2 + 7\] that one works too so the solution is in fact (-2,5)

OpenStudy (anonymous):

that system has 1 solution, is consistent, and they are independent

OpenStudy (darkbluechocobo):

o: Thank you sir again saved me on that

OpenStudy (anonymous):

No problem!

OpenStudy (darkbluechocobo):

Theoretical High five I got an 80

OpenStudy (darkbluechocobo):

which mean my average is still a 93 Whoop WHoop

OpenStudy (darkbluechocobo):

Thank you so so so so sso so so so much

OpenStudy (anonymous):

haha very nice!

OpenStudy (darkbluechocobo):

May I add I hate my algebra teacher

OpenStudy (anonymous):

Been there, done that. Just gotta try and teach yourself if the teacher isnt doing a good job or ask a ton of questions.